Tom's home is 1,800 m from his school. One morning he walked part of the way to school and ran the rest. If it took him 20 minutes or less to get to school, and he walks at 70 m/min and runs 210 m/min, how far did he run?
Answer by lwsshak3(6763)

Tom's home is 1,800 m from his school. One morning he walked part of the way to school and ran the rest. If it took him 20 minutes or less to get to school, and he walks at 70 m/min and runs 210 m/min, how far did he run?
**
let x=distance Tom ran
1800-x=distance Tom walked
travel time=distance/rate of speed
..
x/210+(1800-x)/70=20
LCD:210
x+3(1800-x)=4200
x+5400-3x=4200
2x=1200
x=600 m
ans:
distance Tom ran=600 m